what would be the electric field due to 
Non-Conducting Solid Sphere (Gaussian surgace)
tommorow is my school fa

Grade:10

2 Answers

Abhinav
8 years ago
(i) Outside point

E=\frac{1}{4\pi \epsilon _{o}}\frac{q}{r^{2}}

(ii) Inside point q = total charge
\rho= volume density of charge

E=\frac{1}{4\pi \epsilon_{o}}\frac{qr}{\rho ^{3}}
grenade
8 years ago
u can solve this question by assuming that sphere to be a gaussian surface 
the then u will get the answer as stated above